A risk-free investment of $6000 will return 3.5%. A risky $6000 investment has a 21% chance of defaulting and returning only $3500. How much must the risky investment promise to return?

            

Group of answer choices

15.5%

9.5%

10.5%

11.8%                                                 
Answer #1

As Risk Free Return is 3.5%,

Expected Return of Risky Investment should be at least 3.5%

Therefore, Return promised at which, expected return will be 3.5%, is the return that it MUST promise.

Expected Return = (Probability of Return×Promised Return)-(Probability of Loss×Possible Loss)

Expected Return = 3.5%

Probability of Loss = 0.21 Therefore, Probability of Return = 1-0.21 = 0.79

Possible Loss = (6000-3500)/6000 = 41.6667

Applying in the formula,

3.5 = (0.79×Promised Return)-(0.21×41.6667)

3.5 + 8.75 = 0.79×Promised Return

Therefore, Promised Return = 12.25/0.79 = 15.5%

(If this was helpful then please rate positively. Thank You:)